The price of a stock is down 5% today. What percent increase tomorrow would return
the stock to its original price?

Answer by Alan3354(69443)

The price of a stock is down 5% today. What percent increase tomorrow would return the stock to its original price?
---------
p = starting price
Down 5% --> 0.95p
Using 0.95p as the base, increasing to 1.0p is an increase of (1 - 0.95)/0.95
= 5/95 = 1/19
=~ 0.05263 = 5.26%